Write an equation of the line containing the given point and perpendicular to the given line.
(0,9) 4x+7y=2

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Write an equation of the line containing the given point and perpendicular to the given line.
(0,9) 4x+7y=2
---
Find the slope of the given line:
y = (-4/7)x+(2/7)
slope = -4/7
----------------------
Any perpendicular line must have slope = 7/4
-------
y-intercept is 9
-----
Required equation is y = (7/4)x+9
=====================================
Cheers,
Stan H.
===========

Answer by ewatrrr(24785)

 
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
line containing point (0,9) and perpendicular to 4x+7y=2
4x+7y=2 OR y = (-4/7)x + 2/7 slope = (-4/7)
New Line: slope is the negative reciprocal of (-4/7)
y = (7/4)x + b |b is the y-intercept (0,9) 9 = b
y = (7/4)x + 9